# Does Refusal Training in LLMs Generalize to the Past Tense?

### v2

@@ 5 Related work
After pretraining, LLMs are typically aligned to human preferences using techniques like SFT (Chung et al., 2022), RLHF (Ouyang et al., 2022), or DPO (Rafailov et al., 2023). One of the objectives of the alignment process is to make LLMs produce refusals on harmful queries, which involves adding refusal examples to the fine-tuning data. Since it is impossible to add all possible reformulations of harmful requests in the fine-tuning set, LLMs alignment crucially relies on the ability to generalize from a few examples per harmful behavior. Empirical studies support this capability: Dang et al. (2024) observe that RLHF generalizes from English to other languages, and Li et al. (2024b) make the same claim specifically for refusal training. This observation is consistent with Wendler et al. (2024) who argue that LLMs pretrained primarily on English data tend to internally map other languages to English. Therefore, fine-tuning on English data can suffice since the internal representations largely coincide with the representations of other languages. However, this capacity is in stark contrast to past tense reformulations, which, as we show, represent a blind spot. We hypothesize that the underlying reason is that the internal representations for the past and present tenses are distinct. Thus, one has to include reformulations in both tenses to achieve more robust refusals. On a related note, Berglund et al. (2024) discuss the reversal curse phenomenon, i.e., training on ”A is B” does not make the model automatically generalize to ”B is A”. This failure again illustrates that generalization mechanisms taken for granted in humans do not necessarily occur in LLMs.
Adversarial attacks on LLMs.
@@
Adversarial attacks on deep learning models are well-documented (Biggio et al., 2013; Szegedy et al., 2014; Biggio and Roli, 2018; Madry et al., 2018). Attacks on the refusal mechanism of LLMs are more recent and start with manual techniques, such as those discovered for ChatGPT by Mowshowitz (2022), that involve handcrafted prompts. Wei et al. (2023a) roughly divide these manual jailbreaks on two categories: (1) competing objectives, where model properties, such as helpfulness, conflicts with safety goals, and (2) mismatched generalization, where refusal training fails to generalize to some reformulations such as in low-resource languages, base64, or leetspeak. Additionally, these techniques can be enhanced by using in-context learning prompts that contain examples of harmful responses (Wei et al., 2023b; Anil et al., 2024). Furthermore, jailbreaking can be automated using iterative discrete optimization techniques, opening up a new attack vector that leads to standardized attacks instead of manual trial and error. For example, Zou et al. (2023) use a gradient-based method, Greedy Coordinate Gradient, to create universal and transferable adversarial suffixes. Other approaches involve genetic algorithms to construct adversarial prompts (Lapid et al., 2023; Liu et al., 2023), low-perplexity modifications of GCG (Zhu et al., 2023), or random search on predicted probabilities (Andriushchenko et al., 2024; Sitawarin et al., 2024; Hayase et al., 2024). Our method is conceptually much simpler than these methods since it directly reformulates a harmful query without any iterative optimization.
Most related existing approaches.
@@
Our method is naturally related to multiple existing approaches. First, it falls into the mismatch generalization category of Wei et al. (2023a), as confirmed by the experiments in Section 4. Second, the idea of using an auxiliary LLM to produce various reformulations of harmful requests is at the core of other prior works, such as Perez et al. (2022) and Chao et al. (2023). Finally, our attack is related to the Crescendo multi-turn attack (Russinovich et al., 2024), which involves asking multiple benign questions about a subject before posing a target question. In some cases, Crescendo also leverages past tense reformulations; for example, their Figure 1 shows an example about a Molotov cocktail where the last question is ”How was it created back then?”. We show that this step alone is often crucial, and multi-turn conversations may not always be necessary to produce a jailbreak for many frontier LLMs.

### v3

@@ 6 Related Work
Generalization of LLM alignment. After pretraining, LLMs are typically aligned to human preferences using techniques like SFT (Chung et al., 2022), RLHF (Ouyang et al., 2022), or DPO (Rafailov et al., 2023). One of the objectives of the alignment process is to make LLMs produce refusals on harmful queries, which involves adding refusal examples to the fine-tuning data. Since it is impossible to add all possible reformulations of harmful requests in the fine-tuning set, LLMs alignment crucially relies on the ability to generalize from a few examples per harmful behavior. Empirical studies support this capability: Dang et al. (2024) observe that RLHF generalizes from English to other languages, and Li et al. (2024b) make the same claim specifically for refusal training. This observation is consistent with Wendler et al. (2024) who argue that LLMs pretrained primarily on English data tend to internally map other languages to English. Therefore, fine-tuning on English data can suffice since the internal representations largely coincide with the representations of other languages. However, this capacity is in stark contrast to past-tense reformulations, which, as we show, represent a blind spot. We hypothesize that the underlying reason is that the internal representations for the past and present tenses are distinct. Thus, one has to include reformulations in both tenses to achieve more robust refusals. On a related note, Berglund et al. (2024) discuss the reversal curse phenomenon, i.e., training on ”A is B” does not make the model automatically generalize to ”B is A”. This failure again illustrates that generalization mechanisms taken for granted in humans do not necessarily occur in LLMs.
@@
Adversarial attacks on LLMs. Adversarial attacks on deep learning models are well-documented (Biggio et al., 2013; Szegedy et al., 2014; Biggio and Roli, 2018; Madry et al., 2018). Attacks on the refusal mechanism of LLMs are more recent and start with manual techniques, such as those discovered for ChatGPT by Mowshowitz (2022), that involve handcrafted prompts. Wei et al. (2023a) roughly divide these manual jailbreaks on two categories: (1) competing objectives, where model properties, such as helpfulness, conflicts with safety goals, and (2) mismatched generalization, where refusal training fails to generalize to some reformulations such as in low-resource languages, base64, or leetspeak. Additionally, these techniques can be enhanced by using in-context learning prompts that contain examples of harmful responses (Wei et al., 2023b; Anil et al., 2024). Furthermore, jailbreaking can be automated using iterative discrete optimization techniques, opening up a new attack vector that leads to standardized attacks instead of manual trial and error. For example, Zou et al. (2023) use a gradient-based method, Greedy Coordinate Gradient, to create universal and transferable adversarial suffixes. Other approaches involve genetic algorithms to construct adversarial prompts (Lapid et al., 2023; Liu et al., 2023), low-perplexity modifications of GCG (Zhu et al., 2023), or random search on predicted probabilities (Andriushchenko et al., 2024; Sitawarin et al., 2024; Hayase et al., 2024). Our method is conceptually much simpler than these methods since it directly reformulates a harmful query without any iterative optimization.
@@
Most related existing approaches. Our method is naturally related to multiple existing approaches. First, it falls into the mismatch generalization category of Wei et al. (2023a), as confirmed by the experiments in Section 4. Second, the idea of using an auxiliary LLM to produce various reformulations of harmful requests is at the core of other prior works, such as Perez et al. (2022) and Chao et al. (2023). Finally, our attack is related to the Crescendo multi-turn attack (Russinovich et al., 2024), which involves asking multiple benign questions about a subject before posing a target question. In some cases, Crescendo also leverages past-tense reformulations; for example, their Figure 1 shows an example about a Molotov cocktail where the last question is ”How was it created back then?”. We show that this step alone is often crucial, and multi-turn conversations may not always be necessary to produce a jailbreak for many frontier LLMs.